OpenStudy (anonymous):

Mark currently has a balance of $973.70 in an account he has held for 17 years. He opened the account with an initial deposit of $700. What is the simple interest on the account?

OpenStudy (anonymous):

16.1

OpenStudy (anonymous):

what do i do after that

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

16.1 is not an option

OpenStudy (anonymous):

8.2?

OpenStudy (anonymous):

I=P r T 973.70 = 700 x R x17 r = 973.70/700/17

OpenStudy (anonymous):

that should give you the percentage sorry my bad

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what do i do

OpenStudy (anonymous):

When you know the principal amount, the initial amount and the time. The amount of interest can be calculated by using the formula:I = Prt which can be moved around algebraically r= I/P/t your variables are I = 973.70 P = 700 t= 17 years